# Quantifying resilience for distribution system customers with SALEDI

### v1

@@ I Introduction
There has been a proliferation of resilience concepts and suggested metrics [11, 19, 17], including advances describing and improving distribution system resilience [9, 20, 22]. However, tracking distribution system resilience with quantitative indices calculated from utility outage data, similarly to tracking reliability with the System Average Interruption Duration Index SAIDI, has remained elusive. In this paper we develop a novel System Average Large Event Duration Index SALEDI to quantify distribution system resilience for customers in terms of customer minutes interrupted in large events.
@@
Resilience addresses rare, high-impact extreme events, and these are inherently challenging to track with metrics because of their rarity and the high variability in the event magnitude. For example, it is well known that the index SAIDI, that describes annual reliability for distribution customers with CMIp, is erratic if major event days are not excluded from the calculation [25, 6]. More recently, experience with recorded distribution utility data shows that large event customer costs and customer minutes interrupted are highly variable over orders of magnitude [2, 3, 18]. The high variability is due to heavy tails in the distribution of event magnitudes. The heavy tails and high variability do seem to be an inherent feature of extreme events in power systems [^1] Power interrupted in transmission system blackouts is also well known to be heavy-tailed and highly variable over orders of magnitude [4, 13]..
One consequence of the high variability is that some seemingly straightforward metrics calculated from real utility data that include extreme events are so erratic as to be impractical. The impracticality manifests as these metrics requiring impractically large amounts of data to give a value that is not unduly fluctuating. To grapple with this problem, it is essential to check the statistical variability of any proposed resilience metrics, and this, while addressed in our recent initial work in [2, 3], has not seemed to be addressed at all in the rest of the literature on distribution system resilience. Accordingly, when this paper proposes SALEDI, we quantify its variability and demonstrate its practical computation with only a few years of large event utility data.
@@
  • Interprets SALEDI and ALED as areas under the tail of exceedance curves and relates them to logarithmic versions of Conditional Value at Risk CVAR. Interprets SAIDI as the area under an exceedance curve to better link current industry practice with risk analysis methods.
@@ I-A Literature review
A significant advance towards directly quantifying power system resilience is using Conditional Value at Risk (CVAR) [^2] Alternative names for CVAR are average value-at-risk, expected shortfall, tail conditional expectation, mean expected loss, mean shortfall, and tail VAR. to describe the risk of large events [24, 22, 19]. CVAR is the mean value of the large events, where the value can be energy not served or other measures. The importance of CVAR is that it directly describes the resilience risk of a top percentage of large events. Moreover, the risk of large events measured by CVAR can be balanced with the reliability risk of smaller events by optimizing a linear combination of these risks. However, as shown in [2] for some distribution utility customer cost data and as suggested in [4] for transmission utility data, CVAR can be impractical to calculate from observed data because the high variability of blackout cost requires too many samples to get a reliable estimate. The SALEDI metric of this paper shows how to solve this problem for CMIp with a logarithmic transformation.
@@
Heavy tails in financial data are addressed in [21], focusing mainly on heavy tails with exponent $\alpha>1$. A power law tail is approximated with a Hill estimator and used to estimate value at risk in [21, Prop. 8.13].
There is some previous work using a logarithmic transformation for resilience metrics. Pandey et al. [18] considers the event customer hours divided by the number of customers interrupted and then takes a logarithm to define an Area Index of Resilience metric AIR. A Restoration Efficiency metric RE is the logarithm of the number of repair crews divided by the number of outages. Adding RE and AIR gives a logarithmic REPAIR metric that combines both customer impact and repair efficiency. The logarithm is justified for these metrics by the quantities varying over orders of magnitude for different events.
@@ II-A SAIDI including major event days is highly variable
When attempting to use SAIDI with major event days as a metric, its value fluctuates significantly from year to year, as shown in Fig. 1. It is not credible that the reliability or the resilience is changing drastically each year; SAIDI with major event days is responding to the excessive variability in the CMIp of large events that occur in each year rather than the utility’s reliability or resilience performance. After all, the distribution system reliability and resilience vary relatively slowly over many years as upgrades and changes in equipment and procedures are made, and do not dramatically decrease when the system is hit by a rare extreme of weather causing a very large blackout. We need sufficient stability in the reliability and resilience metrics of power system performance to guide mitigation and investment decisions.
@@ II-C Event customer minutes interrupted and large events
For resilience we focus on large events by setting a threshold $M_{\rm large}$ so that the large events are the events with CMIp $M\geq M_{\rm large}$. Section V-B explains how to choose a suitable value of $M_{\rm large}$.
@@ II-D SALEDI definition
Note the simplicity of SALEDI: First add up over a period of $n_{\rm year}$ years the logarithm of the CMIp in the large events divided by $M_{\rm large}$. Then scale to an annual index by dividing by $n_{\rm year}$. SALEDI is a system-average metric that assesses impact per customer served, because $m_{i}$ and $M_{i}$ are normalized by the number of customers served in (1).
II-E SALEDI incorporates frequency and magnitude of large events
@@
$\text{where }\qquad f_{\rm large}$ $=\frac{n_{\rm large}}{n_{\rm year}}$ (6)
is the annual frequency of large events estimated over the period of $n_{\rm year}$ years, and the Average Large Event Duration metric
@@ II-F Results
Table I describes the utility data, calculated quantities, and the values of SALEDI and ALED for the complete data of the last $n_{\rm year}$ years of each utility, where $n_{\rm year}$ is determined to ensure an $\rm RSE_{SAL}\approx 0.1$ as discussed in section V. Fig. 2 shows how SALEDI varies when calculated over a sliding window of length $n_{\rm year}$. Indeed, Fig. 2 shows SALEDI describing the power system resilience with respect to large event customer impacts with low variability. The low variability of SALEDI is explained and quantified in section III. The only heavily urbanized case we consider is utility 5, and this might explain its lower SALEDI and ALED and higher resilience.
@@
$f_{\rm large}^{\rm all}$ 133.6 73.5 65.5 57.1 40.4
$\overline{n}_{\rm large}$ 267 220 197 171 202
$n_{\rm year}$ 2 3 3 3 5
$\rm RSE_{SAL}$ 0.09 0.10 0.10 0.11 0.10
${\rm RSE}_{{\rm Pb}}$ 40.06 60.96 25.56 33.68 13.44
${\rm RSE}_{\rm LNb}$ 20.09 64.93 21.44 8.74 2.66
$n_{\rm large}$ 290 150 203 168 168
$f_{\rm large}$ 145.0 50.0 67.7 56.0 33.6
SALEDI 157.9 68.3 98.6 52.5 22.0
ALED 1.09 1.37 1.46 0.94 0.65
Figure 2: Tracking SALEDI over years. For each utility, a sliding window of duration $n_{\rm year}$ is used.
@@ III Technical basis of SALEDI
In this section, we develop and explain the theory behind SALEDI, calculate its statistical variation, and relate it to the area under an exceedance function.
@@ III-A Observed exceedance functions
where $I$ is the indicator function. $\overline{F}(M)$ describes risk in the sense that it combines event frequency and event impact in terms of CMIp. Fig. 3 shows the exceedance functions $\overline{F}(M)$ for five utilities on a log-log plot. For quantifying resilience, we are interested in characterizing the tail $M\geq M_{\rm large}$ of the exceedance function. We choose an $M_{\rm large}$ such that the tail $M\geq M_{\rm large}$ has an approximate straight-line behavior on the log-log plot. The slope magnitude of the tail is denoted by $\alpha$. $M_{\rm large}$ and $\alpha$ are shown in Table I. More discussion on the selection of $M_{\rm large}$ is in section V-B.
Figure 3: Exceedance functions $\overline{F}(M)$ for all events of utilities 1-5; log-log plot
@@ III-B Linear approximation to the exceedance function tail
Figure 4: Tails of the exceedance functions of Fig. 3 with normalized event CMIp $P=M/M_{\rm large}$ on the horizontal axis with a log scale.
@@
This follows since $\overline{F}_{X}(x)={\rm P}[X>x]={\rm P}[\ln P>x]={\rm P}[P>e^{x}]=\overline{F}_{P}[e^{x}]=e^{-\alpha x}$. That is, the logarithm of the heavy-tailed Pareto data with slope magnitude $\alpha$ has a light-tailed exponential distribution with rate $\alpha$; also see Fig. 5.
@@
The linear approximation of the power-law tail on a log-log plot has slope magnitude $\alpha$. Equation (7) shows that $\rm ALED$} estimates the mean of the exponential distribution (10), which is $\alpha^{-1}$. Therefore ${\rm ALED}^{-1}$ estimates the slope magnitude $\alpha$ of the power law tail. Indeed ${\rm ALED}^{-1}$ is the Hill estimator [^6] Determining a suitable value of $M_{\rm large}$ is a well-known delicate issue for the Hill estimator. for $\alpha$ [16, 23]. This important relation between ALED and the slope magnitude $\alpha$ shows how ALED describes the power law tail of direct interest to resilience risk. Moreover, the power law slope magnitude captures the linear trend of the large event CMIp, and this trend governs the largest blackouts that may be experienced as discussed in subsection IV-D. Since $\rm ALED$} is a factor of $\rm SALEDI$} as shown in (5), $\rm SALEDI$} also incorporates the linear trend of the large event CMIp.
@@ III-D Statistical variation of SALEDI
To estimate the statistical variation of the number of large events, let $\overline{n}_{\rm large}$ be the average number of large events in $n_{\rm year}$ years of data (see (37)). Then the number of large events $n_{\rm large}$ for $n_{\rm year}$ years of data is modeled as a Poisson[$\overline{n}_{\rm large}$] random variable.
@@
[…]
${\rm RSE}_{\rm SAL}=\frac{\sqrt{1+{\rm RSE}[X]^{2}}}{\sqrt{\overline{n}_{\rm large}}}=\frac{\sqrt{2}}{\sqrt{\overline{n}_{\rm large}}}$ (12)
${\rm RSE}_{\rm ALED}=\frac{{\rm RSE}[X]}{\sqrt{\overline{n}_{\rm large}}}=\frac{1}{\sqrt{\overline{n}_{\rm large}}}$ (13)
@@ III-E1 Area interpretation for SAIDI
Let $m_{1},m_{2},...,m_{k}$ be the CMIp in each outage including major event days, and consider the frequency exceedance function
@@
Then, using (15), SAIDI including major event days is the area under the exceedance function:
@@
Also $\overline{G}(0)=k=$ number of outages. There is a similar interpretation of SAIDI without major event days using an exceedance function for the outage CMIp excluding major event days.
III-E2 Area interpretation for SALEDI
Let $M_{1}$$M_{2}$$M_{n_{\rm large}}$ be the CMIp for large events, and consider the annual frequency exceedance function of this tail data
$\overline{G}(L)$ $=\frac{1}{n_{\rm year}}\sum_{i=1}^{n_{\rm large}}I[\ln\frac{M_{i}}{M_{\rm large}}>L]$ (18)
Then, using (15), SALEDI is the area under $\overline{G}$:
$\int_{0}^{\infty}$ $\!\!\overline{G}(L)dL=\frac{1}{n_{\rm year}}\sum_{i=1}^{n_{\rm large}}\ln\frac{M_{i}}{M_{\rm large}}={\rm SALEDI}$ (19)
Moreover, $\overline{G}(0)=n_{\rm large}/{n_{\rm year}}=f_{\rm large}$.
Similarly, consider the probability exceedance function
$\overline{F}(L)$ $=\frac{1}{n_{\rm large}}\sum_{i=1}^{n_{\rm large}}I[\ln\frac{M_{i}}{M_{\rm large}}>L]$ (20)
Then, using (16) and as shown in Fig. 5(c), ALED is the area under $\overline{F}$:
$\int_{0}^{\infty}$ $\!\!\overline{F}(L)dL=\frac{1}{n_{\rm large}}\sum_{i=1}^{n_{\rm large}}\ln\frac{M_{i}}{M_{\rm large}}={\rm ALED}$ (21)
In summary, if the frequency exceedance curve variable is outage CMIp then the area under the curve is SAIDI, and if the variable is the log of normalized large event CMIp scaled by $1/n_{year}$, then the area under this tail curve is SALEDI, as shown in Fig. 5(d).
@@ IV Comparing SALEDI with alternatives
This section explains 1) why the logarithm in SALEDI and ALED is necessary by describing the high variability when the logarithm is removed, 2) compares SALEDI and ALED with SAIDI and CVAR, and 3) discusses how ALED describes the trend governing the largest blackouts that may not have yet occurred.

### v2

# Author preprint March 2026 Quantifying resilience for distribution system customers with SALEDI

@@ I Introduction
There has been a proliferation of resilience concepts and suggested metrics [11, 18, 16], including advances describing and improving distribution system resilience [9, 19, 21]. However, tracking distribution system resilience with quantitative indices calculated from utility outage data, similarly to tracking reliability with the System Average Interruption Duration Index SAIDI, has remained elusive. In this paper we develop a novel System Average Large Event Duration Index SALEDI to quantify distribution system resilience in terms of customer minutes interrupted in large events.
@@
Resilience addresses rare, high-impact extreme events, and these are inherently challenging to track with metrics because of their rarity and the high variability in the event magnitude. For example, it is well known that SAIDI, which describes annual reliability for distribution customers with CMIp, is erratic if major event days are not excluded from the calculation [24, 6]. More recently, experience with recorded distribution utility data shows that large event customer costs and customer minutes interrupted are highly variable over orders of magnitude [2, 3, 17]. The high variability is due to heavy tails in the distribution of event magnitudes. The heavy tails and high variability do seem to be an inherent feature of extreme events in power systems [^1] Power interrupted in transmission system blackouts is also well known to be heavy-tailed and highly variable over orders of magnitude [4]..
One consequence of the high variability is that some seemingly straightforward metrics calculated from real utility data that include extreme events are so erratic as to be impractical. The impracticality manifests as these metrics requiring impractically large amounts of data to give a value that is not unduly fluctuating. To grapple with this problem, it is essential to check the statistical variability of any proposed resilience metrics, and this, while addressed in our recent initial work in [2, 3], has not seemed to be addressed at all in the rest of the literature on distribution system resilience. Accordingly, when this paper proposes SALEDI, we quantify its variability and demonstrate that it can be practically computed using only a few years of large-event utility data.
@@
  • Interprets SALEDI and ALED as areas under the tail of exceedance curves and relates them to logarithmic versions of Conditional Value at Risk CVAR. Interprets SAIDI as the area under an exceedance curve to better link risk analysis methods with current industry practice.
@@ I-A Literature review
A significant advance towards directly quantifying power system resilience is using Conditional Value at Risk (CVAR) [^2] Alternative names for CVAR are average value-at-risk, expected shortfall, tail conditional expectation, mean expected loss, mean shortfall, and tail VAR. to describe the risk of large events [23, 21, 18]. CVAR is the mean value of the large events, where the value can be energy not served or other measures. The importance of CVAR is that it directly describes the resilience risk of a top percentage of large events. Moreover, the risk of large events measured by CVAR can be balanced with the reliability risk of smaller events by optimizing a linear combination of these risks. However, as shown in [2] for some distribution utility customer cost data and as suggested in [4] for transmission utility data, CVAR can be impractical to calculate from observed data because the high variability of blackout cost requires too many samples to get a reliable estimate. The ALED metric of this paper shows how to solve this problem for CMIp with a logarithmic transformation.
@@
Heavy tails in financial data are addressed in [20], focusing mainly on heavy tails with exponent $\alpha>1$. A power law tail is approximated with a Hill estimator and used to estimate value at risk in [20, Prop. 8.13].
There is some previous work using a logarithmic transformation for resilience metrics. Pandey [17] considers the event customer hours divided by the number of customers interrupted and then takes a logarithm to define an Area Index of Resilience metric AIR. A Restoration Efficiency metric RE is the logarithm of the number of repair crews divided by the number of outages. Adding RE and AIR gives a logarithmic REPAIR metric that combines both customer impact and repair efficiency. The logarithm is justified for these metrics by the quantities varying over orders of magnitude for different events.
@@ II-A SAIDI including major event days is highly variable
When attempting to use SAIDI with major event days as a metric, its value often fluctuates significantly from year to year, as shown in Fig. 1. The dominant source of fluctuation is the random occurrence and variable magnitude of large events. That is, SAIDI with major event days is responding more to the very high variability in the large events that can occur in each year than to the power system’s underlying reliability or resilience. After all, the distribution system reliability and resilience vary relatively slowly over many years as upgrades and changes in equipment and procedures are made. We need resilience metrics to more stably track power system performance as it responds to weather extremes in order to guide mitigation and investment decisions.
@@ II-C Event customer minutes interrupted and large events
For resilience we focus on large events by setting a threshold $M_{\rm large}$ so that the large events are the events with CMIp $M\geq M_{\rm large}$. Section V-A explains how to choose a suitable value of $M_{\rm large}$.
@@ II-D SALEDI definition
Note the simplicity of SALEDI: First add up over a period of $n_{\rm year}$ years the logarithm of the CMIp in the large events normalized by $M_{\rm large}$. Then scale to an annual index by dividing by $n_{\rm year}$. SALEDI is a system-average metric that assesses impact per customer served, because $m_{i}$ and $M_{i}$ are normalized by the number of customers served in (1).
II-E SALEDI incorporates large event frequency & magnitude
@@
$\text{where }\qquad f_{\rm large}$ $=n_{\rm large}/n_{\rm year}$ (6)
is the annual frequency of large events estimated over $n_{\rm year}$ years, and the Average Large Event Duration metric
@@ II-F Results
Table I describes the utility data, calculated quantities, and the values of SALEDI and ALED for the complete data of the last $n_{\rm year}^{\rm min}$ years of each utility, where $n_{\rm year}^{\rm min}$ is determined to ensure that SALEDI’s relative standard error $\rm RSE_{SAL}=0.1$ as discussed in section V. Fig. 2 shows how SALEDI varies when calculated over a sliding window of length $n_{\rm year}^{\rm min}$ rounded to an integer number of years. Indeed, Fig. 2 shows SALEDI describing the power system resilience with respect to large event customer impacts with low variability. The low variability of SALEDI is explained and quantified in section III. The only heavily urbanized case we consider is utility 5, and this might explain its lower SALEDI and ALED and higher resilience.
@@
${\rm RSE}_{{Pb}}$ 40.06 60.96 25.56 33.68 13.44
${\rm RSE}_{LNb}$ 39.48 64.94 21.44 8.74 2.66
$n_{{\rm large}\,Pb}^{\rm minSPL}$ $1.6\mathrm{E}{5}$ $3.7\mathrm{E}{5}$ $6.5\mathrm{E}{4}$ $1.1\mathrm{E}{5}$ $1.8\mathrm{E}{4}$
$n_{{\rm large}LNb}^{\rm minSPL}$ $1.6\mathrm{E}{5}$ $4.2\mathrm{E}{5}$ $4.6\mathrm{E}{4}$ $7700$ $800$
following rows calculated from the last $n_{\rm year}^{\rm min}$ years of data so that
[…]
$f_{\rm large}$ 133.6 73.5 65.5 57.1 40.4
SALEDI 134.8 111. 93.7 55.9 26.7
ALED 1.01 1.51 1.43 0.98 0.66
Figure 2: Tracking SALEDI over years. For each utility, a sliding window of duration $n_{\rm year}^{\rm min}$ rounded to an integer is used.
@@ III Technical basis of SALEDI
This section develops and explains the theory behind SALEDI, calculates its statistical variation, and relates it to the area under an exceedance function.
@@ III-A Observed exceedance functions
where $I$ is the indicator function. $\overline{F}(M)$ describes risk in the sense that it combines event frequency and event impact in terms of CMIp. Fig. 3 shows the exceedance functions $\overline{F}(M)$ for the five utilities on a log-log plot. For quantifying resilience, we are interested in characterizing the tail $M\geq M_{\rm large}$ of the exceedance function. We choose an $M_{\rm large}$ such that the tail $M\geq M_{\rm large}$ has an approximate straight-line behavior on the log-log plot. The slope magnitude of the tail is denoted by $\alpha$. $M_{\rm large}$ and $\alpha$ are shown in Table I. More discussion on the selection of $M_{\rm large}$ is in section V-A.
Figure 3: Exceedance functions $\overline{F}(M)$ for all events of utilities 1-5; log-log plot.
@@ III-B Linear approximation to the exceedance function tail
Figure 4: Tails of the exceedance functions of Fig. 3 with normalized event CMIp $P=M/M_{\rm large}$ on the horizontal axis; log-log plot.
@@
This follows since $\overline{F}_{X}(x)={\rm P}[X>x]={\rm P}[\ln P>x]={\rm P}[P>e^{x}]=\overline{F}_{P}(e^{x})=e^{-\alpha x}$. That is, the logarithm of the heavy-tailed Pareto data with slope magnitude $\alpha$ has a light-tailed exponential distribution with rate $\alpha$; also see Fig. 5.
@@
The linear approximation of the power-law tail on a log-log plot has slope magnitude $\alpha$. Equation (7) shows that $\rm ALED$} estimates the mean of the exponential distribution (10), which is $\alpha^{-1}$. Therefore ${\rm ALED}^{-1}$ estimates the slope magnitude $\alpha$ of the power law tail. Indeed, ${\rm ALED}^{-1}$ is the Hill estimator [^6] Determining a suitable value of $M_{\rm large}$ is a well-known delicate issue for the Hill estimator. for $\alpha$ [15, 22]. This important relation between ALED and the slope magnitude $\alpha$ shows how ALED describes the power law tail of direct interest to resilience risk. Moreover, the power law slope magnitude captures the linear trend of the large event CMIp, and this trend governs the largest blackouts that may be experienced as discussed in subsection IV-D. Since $\rm ALED$} is a factor of $\rm SALEDI$} as shown in (5), $\rm SALEDI$} also incorporates the linear trend of the large event CMIp.
@@ III-D Statistical variation of SALEDI
To estimate the statistical variation of the number of large events, let $N_{\rm large}$ be the random variable that is the number of large events in $n_{\rm year}$ years of data. Then $N_{\rm large}$ is modeled as a Poisson[$n_{\rm large}$] random variable.
@@
Since $X$ follows the exponential distribution (10) [^7] Upper bounding (10) with the maximum normalized CMIp $M_{\rm max}/M_{\rm large}$ considered in subsection IV-A makes no practical difference., ${\rm E}X=\alpha^{-1}$ and ${\rm Var}X=\alpha^{-2}$ so that ${\rm RSE}_{X}=\sigma[X]/{\rm E}X=1$. Then Appendix B calculates the RSE of SALEDI and ALED as
${\rm RSE}_{\rm SAL}=\frac{\sqrt{1+({\rm RSE}_{X})^{2}}}{\sqrt{n_{\rm large}}}=\frac{\sqrt{2}}{\sqrt{n_{\rm large}}}$ (12)
${\rm RSE}_{\rm ALED}=\frac{{\rm RSE}_{X}}{\sqrt{n_{\rm large}}}=\frac{1}{\sqrt{n_{\rm large}}}$ (13)
@@ III-E1 Area interpretation for SAIDI
Let $m_{1},m_{2},...,m_{k}$ be the CMIp in each outage, and consider the frequency exceedance function
@@
Then, using (15), SAIDI is the area under the frequency exceedance function (see example in Fig 6):
@@
Also $\overline{G}(0)=k=$ number of outages.
Figure 6: Simple example of 3 outages with CMIp values 10, 20, 40 respectively. Area under exceedance curve $\overline{G}=10+20+40={\rm SAIDI}$ for the 3 outages.
III-E2 Area interpretation for SALEDI and ALED
Let $M_{1}$$M_{2}$$M_{n_{\rm large}}$ be the CMIp for large events, and consider the annual frequency exceedance function of the tail
$\overline{G}_{S}(L)$ $=\frac{1}{n_{\rm year}}\sum_{i=1}^{n_{\rm large}}I[\ln\frac{M_{i}}{M_{\rm large}}>L]$ (18)
Using (15), SALEDI is the area under $\overline{G}_{S}$ (see Fig. 5(d)):
$\int_{0}^{\infty}$ $\!\!\overline{G}_{S}(L)dL=\frac{1}{n_{\rm year}}\sum_{i=1}^{n_{\rm large}}\ln\frac{M_{i}}{M_{\rm large}}={\rm SALEDI}$ (19)
Moreover, $\overline{G}_{S}(0)=n_{\rm large}/{n_{\rm year}}=f_{\rm large}$.
Similarly, consider the probability exceedance function for the tail $\overline{F}_{S}(L)=\frac{n_{\rm year}}{n_{\rm large}}G_{S}(L)$. Then, using (16) and as shown in Fig. 5(c), ALED is the area under $\overline{F}_{S}$:
$\int_{0}^{\infty}$ $\!\!\overline{F}_{S}(L)dL=\frac{1}{n_{\rm large}}\sum_{i=1}^{n_{\rm large}}\ln\frac{M_{i}}{M_{\rm large}}={\rm ALED}$ (20)
@@ IV Comparing SALEDI with alternatives
This section 1) explains why the logarithm in SALEDI and ALED is necessary by describing the high variability when the logarithm is removed, 2) compares SALEDI and ALED with SAIDI and CVAR, and 3) discusses how ALED describes the trend governing the largest blackouts.